Angie wants to buy her first home in 5 years. She will need $20,000 as a down payment. Angie just recieved some money from a contest gambling and she has deposited it into her savings account, which is earning 6% and compounds semiannually.

What table will you look at to solve this problem?

What column will look at?

What row will you look at?

What is the factor that you see at the meeting point of the row and column?

How much will Pam have to set aside now to have $20,000 for her down payment?
Answer #1

We need to find the amount required to be deposited today

It will be equal to the present value of amount required in future

Hence, the relevant table is Present value of $1 table

Column consist of interest rate

Semi annual rate = 6%/2 =3%

Hence, column will be 3%

Row = 5*2 =10

Amount required to be deposited today = 20,000*0.744

=$14,880